# EdgeOL: Efficient in-situ Online Learning on Edge Devices

### 30 Jan 2024 v1

@@ 5.5. Semi-supervised Learning
Next, we evaluate the ability of our EdgeOL to utilize unlabeled data by applying semi-supervised learning. We choose the common configuration that only 10% of the training data is labeled (zhai2019s4l, ; wang2022np, ). As shown in Table 8, compared to $Immed.$, EdgeOL delivers 1.23% higher accuracy and saves 76% of execution time and 61% of energy, on average. These results demonstrate that EdgeOL works well in semi-supervised learning scenarios. This is because both DAF and SimFreeze are robust to the insufficient labeled data as i) DAF only needs a very small amount of labeled validation data to get the validation accuracy to adjust the fine-tuning frequency and ii) SimFreeze freezes layers by self-representational similarity, which can be acquired without data labels.
@@ 5.6. Compatibility with Quantization
We also evaluate the compatibility of our EdgeOL with quantization technique (stock2020training, ; gupta2015deep, ). We apply 8-bit fixed-point quantization to weights, activations, the gradient of weights, and the gradient of activations. Following the prior works, we compare the accuracy results since the simulated quantization-aware training is used (zhou2016dorefa, ; wang2018training, ; zhu2020towards, ; zhao2021distribution, ). Table 9 shows that EdgeOL outperforms immediate online learning in 32-bit floating-point baselines with an accuracy improvement of 1.95%. Additionally, when employing 8-bit quantization, EdgeOL achieves a 2.19% higher accuracy. These results suggest that EdgeOL’s advantages are maintained when quantization techniques are used, demonstrating compatibility and robustness.
6. Related Work
A number of approaches have been proposed to reduce the computation costs of DNN models, thereby reducing energy and execution time. E2Train (wang2019e2, ) proposes to drop mini-batches randomly, skip layers selectively, and use low-precision back-propagation during training to reduce the computation costs. (DBLP:conf/iclr/YouL0FWCBWL20, ) designs a low-cost method to train the small but critical subnetworks to achieve the same accuracy as the original neural networks. (hu2022lora, ) proposes to use lightweight low-rank matrices to adapt the weights of original models, slightly sacrificing model representational power to reduce the training costs. However, these and most other prior works focus on offline learning. For online learning, some methods are proposed to filter important data for training to minimize the cost (panda2016conditional, ; wu2021enabling, ). These approaches can effectively reduce training costs. Nevertheless, our approach is complementary to them since it focuses on determining the moment to trigger fine-tuning adaptively and freezing layers selectively.

### 16 Mar 2024 v2

@@ 5.5. Semi-supervised Learning
Next, we evaluate the ability of our EdgeOL to utilize unlabeled data by applying semi-supervised learning. We choose the common configuration that only 10% of the training data is labeled (zhai2019s4l, ; wang2022np, ). As shown in Table 7, compared to $Immed.$, EdgeOL delivers 1.36% higher accuracy and saves 43% energy on average. These results demonstrate that EdgeOL works well in semi-supervised learning cases. This is because both DAF and SimFreeze are robust to the insufficient labeled data as i) DAF only needs a very small amount of labeled validation data to get the validation accuracy to adjust the fine-tuning frequency and ii) SimFreeze freezes layers by self-representational similarity, which can be acquired without data labels.
@@ 5.6. Compatibility with Quantization
We also evaluate the compatibility of our EdgeOL with quantization technique (stock2020training, ; gupta2015deep, ). We apply 8-bit fixed-point quantization to weights, activations, the gradient of weights, and the gradient of activations. Following the prior works, we compare the accuracy results since the simulated quantization-aware training is used (zhou2016dorefa, ; wang2018training, ; zhu2020towards, ; zhao2021distribution, ). Table 8 shows that EdgeOL outperforms immediate online learning in 32-bit floating-point baselines with an accuracy improvement of 2.03%. On the other hand, when employing 8-bit quantization, EdgeOL achieves a 2.11% higher accuracy. These results suggest that EdgeOL’s advantages are maintained when quantization techniques are used, demonstrating compatibility and robustness.
6. Related Works and Discussion
A number of approaches have been proposed to reduce the computation costs of DNN models, thereby reducing energy and execution time. E2Train (wang2019e2, ) proposes to drop mini-batches randomly, skip layers selectively, and use low-precision back-propagation during training to reduce the computation costs. (DBLP:conf/iclr/YouL0FWCBWL20, ) designs a low-cost method to train the small but critical subnetworks to achieve the same accuracy as the original neural networks. (hu2022lora, ) proposes to use lightweight low-rank matrices to adapt the weights of original models, slightly sacrificing model representational power to reduce the training costs. However, these and most other prior works focus on offline learning.
For online learning, there are some works proposed to optimize particularly for online video analytics applications (khani2023recl, ; bhardwaj2022ekya, ). Specifically, RECL (khani2023recl, ) maintains a model zoo and uses the streaming-in training data to fine-tune these models, where the most appropriate model will be selected for inference in different scenarios. Ekya (bhardwaj2022ekya, ) strategically schedules the resources among the training and inference workloads of co-running applications to achieve higher inference accuracy. Due to the continuous and regular nature of video streaming in those applications, these works typically divide the online learning process into multiple short windows (e.g., 200 seconds) and conduct online fine-tuning in each window in a fixed-frequency manner. Some other methods are proposed to filter important data for training to minimize the cost (panda2016conditional, ; wu2021enabling, ), which can effectively reduce training costs. Moreover, (ma2023cost, ) presents a system runtime designed to dynamically configure the episodic memory hierarchy (HEM), where HEM is critical for improving the model performance during online learning. This runtime effectively optimizes both accuracy and energy efficiency. Nonetheless, it is important to emphasize that our approach is complementary to these approaches since we focus on determining the moment to trigger fine-tuning adaptively and freezing layers selectively. We will investigate the incorporation of the above methods in our future works.